OpenStudy (anonymous):

two brothers have 60 candies,and shared it equally,d younger brother sold his at the price of 3candy for 1dollar and made 10dollars,d senior sold his at d price of 2candy for one dollar and made 15dollars dats 25dollars in total..d next time dey decided to sell there candies together 5candies for 2dollars and made 24dollars...how come the difference of 1dollar???

OpenStudy (jack1):

hi @balas100 welcome to OS

OpenStudy (jack1):

3 C for $1 = $1/3 candy \[\large \frac 13 x = $10\] so x = 30 2 C for $1 = $1/2 candy \[\large \frac 12 x = $15\] so x = 30 5 C for $2 = $1/2.5 candy \[\large \frac 25 x = $24\] so x = 60 the reason there's a difference in profits between the 2 is the difference in price 30c @ $0.33 each + 30c @ $0.50 each = average price of : \[\large \frac {$25}{60~candy} = \frac {5}{12} \] whereas: \[\large \frac {$24}{60~candy} = \frac {6}{15} = \frac 25 \] similarly, if the older brother had sold all 60 candy at his first price \[\large \frac {$1}{2~candy} \times 60~candy = $30\] if the younger brother had sold all the candy at his first price \[\large \frac {$1}{3~candy} \times 60~candy = $20\] but the fact is they didn't pick a price in the middle of these , which would be: \[\large \frac {$5}{12~candy} \times 60~candy = $25\] instead they picked a random new price to sell at: 2/5

OpenStudy (anonymous):

wow dat was great bt am still finding your explaination a little confusing